Find the inverse of the function.

f(x) = 6x3 - 3
1
Expert's answer
2014-05-12T03:33:07-0400
To find the inverse function we should express the x from this expression
So we get
f(x)=6x3-3
f(x)+3=6x3
x3=(f(x)+3)/6
x=((f(x)+3)/6)1/3
Therefore we get
f-1(y)=((y+3)/6)1/3